OpenStudy (anonymous):

Solve the equation. Show your work. 8x – 12 = 4x + 24

OpenStudy (anonymous):

why do you keep asking for answers to similar questions. You must try them yourself.

OpenStudy (anonymous):

cause i do not know how to really do them

OpenStudy (anonymous):

then learn or ask, don't just get answers. What don't you understand specifically?

OpenStudy (anonymous):

if you had x+1=2, would you know how to do that?

OpenStudy (anonymous):

do you add 1 to the other side?

OpenStudy (anonymous):

do you mean, x+1=2+1? (like is that what you're saying?)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok so what would you do next?

OpenStudy (anonymous):

thats where i get confused

OpenStudy (anonymous):

haha, ok well. let's start simpler then without all the x stuff so 2=2. right? so if this was some sort of equation and i felt like adding 1 to a side: 2+1=2. does that make sense?

OpenStudy (anonymous):

by the way if it doesn't, that's good because it's not supposed to...